OpenStudy (anonymous):

Please help! Reduce cos(30degrees-theta) - cos(30degrees+theta) to a single term. a. sintheta b. sqrt3 sintheta c. -sintheta d. -sqrt3 sintheta

OpenStudy (asnaseer):

use the formula for cos(A+B) and the one for cos(A-B)

OpenStudy (asnaseer):

do you know these formulas?

OpenStudy (anonymous):

Can you please give them to me? :/

OpenStudy (asnaseer):

have a look on this site: http://www.mathsisfun.com/algebra/trigonometric-identities.html scroll down to the section titled: Angle Sum and Difference Identities

OpenStudy (asnaseer):

its near the very bottom of the page

OpenStudy (anonymous):

ughh I tried solving it but I failed: answer is sin theta http://www.wolframalpha.com/input/?i=%28cos+30+cos+theta+%2B+sin+30+sin+theta%29+-+%28cos30+cos+theta+-+sin30+sin+theta%29 If you could explain how to solve it, I will be grateful asnaseer.

OpenStudy (asnaseer):

@JoshDavoll you should not give the answer out. It is better to guide the user to the answer. That way they learn more.

OpenStudy (anonymous):

Well I've been trying to solve it as well thanks to your guidance. Can you please explain though?

OpenStudy (anonymous):

Oh sorry, asnaseer.

OpenStudy (asnaseer):

what have you tried so far @IloveCharlie ?

OpenStudy (anonymous):

@JoshDavoll so you used wolfram to solve?

OpenStudy (asnaseer):

if you list each step then I can try and spot where you may have made a mistake.

OpenStudy (anonymous):

@IloveCharlie - yeah, I broke the problem into cos(x+y) = cosxcosy-sinxsiny, and cos(x-y) = cosxcosy+sinxsiny.

OpenStudy (anonymous):

Yes, same. I used the formula and link asnaseer provided.

OpenStudy (anonymous):

but after breaking them down to those identities, I got stuck :)

OpenStudy (asnaseer):

you are both thinking along the right lines. so what you need to do is to convert this: cos(30degrees-theta) - cos(30degrees+theta) into: cos(x-y) - cos(x+y) and expand this using the formulas

OpenStudy (asnaseer):

here I have let x=30, y=theta

OpenStudy (asnaseer):

@IloveCharlie can you try doing that and let me know what you get?

OpenStudy (anonymous):

Just a sec

OpenStudy (asnaseer):

ok :)

OpenStudy (anonymous):

When I plugged it into my calc I got 0. I plugged it into wolfram as well http://www.wolframalpha.com/input/?i=cos%2830-theta%29+-+cos%2830%2Btheta%29

OpenStudy (asnaseer):

no - do NOT use a calculator - you don't need to use one here - just use algebra

OpenStudy (asnaseer):

we know cos(x+y) = cos(x)cos(y) - sin(x)sin(y) and cos(x-y) = cos(x)cos(y) + sin(x)sin(y) use these expansions to work out what cos(x-y) - cos(x+y) = ?

OpenStudy (asnaseer):

and leave the result in terms of x and y.

OpenStudy (anonymous):

(cos30cos(theta)+sin30sin(theta)) - (cos30cos(theta)-sin30sin(theta)) right?

OpenStudy (asnaseer):

work with x and y please

OpenStudy (anonymous):

oh right

OpenStudy (asnaseer):

we can substitue in the correct values at the very end

OpenStudy (asnaseer):

@IloveCharlie do you understand or do you want me to explain in more detail?

OpenStudy (anonymous):

(cosxcosy+sinxsiny) - (cosxcosy - sinxsiny) right?

OpenStudy (asnaseer):

yes - and please let ILoveCharlie work this out.

OpenStudy (anonymous):

oh okay..sure..

OpenStudy (anonymous):

Oh yes! I understand now! Thanks so much for the help and explanation!

OpenStudy (asnaseer):

so what do you get for the expression cos(x-y) - cos(x+y) = ?

OpenStudy (anonymous):

sintheta! :)

OpenStudy (asnaseer):

:/ I wasn't asking for the final answer, I wanted to know what you got for: cos(x-y) - cos(x+y) = ?

OpenStudy (anonymous):

Well I was able to plug in x and y to the formula and got sintheta. http://www.wolframalpha.com/input/?i=%28cos30costheta%2Bsin30sintheta%29+-+%28cos30costheta+-+sin30sintheta%29

OpenStudy (asnaseer):

remember wolfram will not be with you in the exams

OpenStudy (anonymous):

:/ sorry. I shouldn't have said anything before..

OpenStudy (asnaseer):

np @JoshDavoll - at least you understand why now :)

OpenStudy (anonymous):

asnaseer, can you still walk me through it? :)

OpenStudy (asnaseer):

sure...

OpenStudy (asnaseer):

so, we started with: we know cos(x+y) = cos(x)cos(y) - sin(x)sin(y) and cos(x-y) = cos(x)cos(y) + sin(x)sin(y) use these expansions to work out what cos(x-y) - cos(x+y) = ?

OpenStudy (anonymous):

thanks! so reduces to (cosxcosy+sinxsiny) - (cosxcosy - sinxsiny)

OpenStudy (asnaseer):

this gives us: cos(x-y) - cos(x+y) = [cos(x)cos(y) + sin(x)sin(y)] - [cos(x)cos(y) - sin(x)sin(y)] = cos(x)cos(y) + sin(x)sin(y) - cos(x)cos(y) + sin(x)sin(y) = 2sin(x)sin(y) agreed?

OpenStudy (anonymous):

sorry, how did the operator change in the 2nd line to a '+' ?

OpenStudy (asnaseer):

because: -( a - b ) = -a -(-b) = -a + b

OpenStudy (asnaseer):

two minuses make a plus

OpenStudy (anonymous):

ahh cool, ok I understand, and how did you get to 2sinxsiny?

OpenStudy (asnaseer):

e.g.: -( 5 - 2 ) = -5 + 2 = -3

OpenStudy (anonymous):

^because the - distributes right?

OpenStudy (asnaseer):

cos(x-y) - cos(x+y) = [cos(x)cos(y) + sin(x)sin(y)] - [cos(x)cos(y) - sin(x)sin(y)] = cos(x)cos(y) + sin(x)sin(y) - cos(x)cos(y) + sin(x)sin(y) = cos(x)cos(y) - cos(x)cos(y) + sin(x)sin(y) + sin(x)sin(y) = 2sin(x)sin(y)

OpenStudy (anonymous):

Ahh! right, I understand all of that.

OpenStudy (anonymous):

^commutative law right?

OpenStudy (asnaseer):

good, so finally if we now plug in the actual values we get:\[\cos(30-\theta)-\cos(30+\theta)=2\sin(30)\sin(\theta)\]

OpenStudy (asnaseer):

(yes commutative law)

OpenStudy (asnaseer):

and we know sin(30) = 0.5

OpenStudy (asnaseer):

therefore:\[\cos(30-\theta)-\cos(30+\theta)=2\sin(30)\sin(\theta)=2\times 0.5\times\sin(\theta)=\sin(\theta)\]

OpenStudy (anonymous):

:) Awesome explaining. Thank you very much, asnaseer.

OpenStudy (asnaseer):

np - and I hope you were also following along IloveCharlie :)

OpenStudy (asnaseer):

it may seem hard at first trying to learn it this way but the benefits will be enormous when it comes to exam time :)

OpenStudy (anonymous):

Yes, I have been. Thank you so much both of you!

OpenStudy (asnaseer):

yw :)